Jefferson Lab HALL A Logbook for July 10 - All EXCEPT Automatic Entries

  • Main INDEX

    Friday

  • 331669 07/30/10 15:07 R. Michaels Qweak helicity into datastream. Looks good.
  • Wednesday

    
    

    Tuesday

  • 331658 07/27/10 14:10 Jack (Techs) Danfysik resets
  • 331657 07/27/10 13:52 Jack (Techs) Argon gas system leak repaired
  • Monday

  • 331656 07/26/10 19:30 ejensen Updated momentum settings in run database
  • Sunday

  • 331655 07/25/10 20:32 Jack (Techs) Argon gas system pressure regulator leaks
  • Friday

  • 331654 07/23/10 20:52 folts target
  • Thursday

  • 331653 07/22/10 12:13 ZAFAR AHMED Correction: Change of pdestal for CR sum in PAW analysis for Moller
  • 331652 07/22/10 12:08 ZAFAR AHMED Change of pdestal for CR sum in PAW analysis for Moller
  • Wednesday

  • 331651 07/21/10 16:05 J. Butler(te Left Dipole
  • 331650 07/21/10 15:45 A.Shahinyan, LHRS VDC test
  • 331647 07/21/10 15:11 A. Shahinyan VDC readout problem
  • 331646 07/21/10 15:10 fansler LQ2
  • Tuesday

  • 331641 07/20/10 16:16 J. Butler (T Installed Phase Current Monitoring System.
  • 331640 07/20/10 16:14 J. Butler (T Replace OPTO 22 Relays
  • 331639 07/20/10 16:12 folts target chamber /tech
  • 331638 07/20/10 16:11 J. Butler (T Replace Phase Balance Relays
  • 331637 07/20/10 15:29 fansler Left arm liquid levels
  • 331636 07/20/10 15:27 fsnsler Right arm CAMAC
  • 331635 07/20/10 13:52 fansler Left arm magnet temps
  • Friday

  • 331634 07/16/10 19:02 Jack (Techs) Right Q2 crossed wiring is internal to the magnet
  • Wednesday

  • 331633 07/14/10 14:16 Fansler RQ1 on warm return
  • Tuesday

  • 331618 07/13/10 15:24 fansler Mass flow controllers
  • 331610 07/13/10 09:38 folts post run checklist
  • 331609 07/13/10 08:23 folts magnets ramped down
  • Monday

  • 331587 07/12/10 08:23 riordan Septum Off
  • 331586 07/12/10 08:08 riordan VDC HV Off
  • 331584 07/12/10 08:03 riordan Owl Shift Summary
  • 331583 07/12/10 08:03 riordan End APEX Test Run, called for restricted access
  • 331581 07/12/10 07:39 riordan Restart CODA
  • 331576 07/12/10 06:46 riordan spot++, run 2081
  • 331573 07/12/10 06:43 rouven VDC check-out and High Rate Coincidence Studies
  • 331572 07/12/10 06:37 riordan Raster 4x6 MCC
  • 331571 07/12/10 06:37 riordan Moving to Ta
  • 331561 07/12/10 06:05 rouven Production of e-, pi+, Boron off Carbon Target
  • 331560 07/12/10 06:01 riordan spot++, 2075
  • 331554 07/12/10 05:53 Neil G-W Spot++ for 2073
  • 331547 07/12/10 05:35 riordan Raster to 12x12 MCC
  • 331546 07/12/10 05:34 riordan Moving to Pb/D #1
  • 331542 07/12/10 04:45 Neil G-W Spot++ for run 2069
  • 331540 07/12/10 04:31 Neil G-W Switch to thick carbon target
  • 331539 07/12/10 04:29 Neil G-W Spot++ for Checkout run 2068
  • 331538 07/12/10 04:28 riordan BeO spot good
  • 331537 07/12/10 04:27 riordan Starting Beam Checkout
  • 331524 07/12/10 01:14 Neil G-W TO screens
  • Sunday

  • 331523 07/11/10 23:22 dalton Swing shift summary
  • 331522 07/11/10 23:16 dalton CHL tripped
  • 331521 07/11/10 23:14 dalton Argon bottle changed
  • 331519 07/11/10 23:05 siqbal spot++, run #2066
  • 331518 07/11/10 22:51 siqbal spot++, run #2065
  • 331514 07/11/10 22:16 dalton Argon gas low pressure
  • 331512 07/11/10 21:53 dalton Beam down ~ 30 mins
  • 331509 07/11/10 21:21 siqbal Spot++, run #2063
  • 331506 07/11/10 21:03 dalton Hall A at 150 uA
  • 331505 07/11/10 20:34 siqbal Spot++, run #2062
  • 331502 07/11/10 19:48 siqbal Spot++, run #2061
  • 331501 07/11/10 19:40 siqbal Spot++, run #2060
  • 331488 07/11/10 17:58 siqbal Spot++ for run #2058
  • 331485 07/11/10 16:57 siqbal Spot++ for run #2057
  • 331482 07/11/10 16:33 siqbal Spot++ for run #2056
  • 331479 07/11/10 16:07 kelleher Day Shift Summary -- June 11, 2010
  • 331477 07/11/10 15:37 boyce spot++ on Run 2055
  • 331474 07/11/10 14:57 boyce spot++ on Run 2054
  • 331471 07/11/10 14:24 boyce spot++ on run 2053
  • 331469 07/11/10 13:46 boyce spot++ on Run 2052
  • 331466 07/11/10 13:05 boyce spot++ on Run 2051
  • 331463 07/11/10 12:18 boyce spot++ on Run 2050
  • 331460 07/11/10 11:33 boyce spot++ on Run 2049
  • 331450 07/11/10 11:05 kelleher back to 120uA
  • 331449 07/11/10 10:57 kelleher beam visible at dump viewer
  • 331448 07/11/10 10:39 boyce spot++ on run 2048
  • 331445 07/11/10 10:03 boyce spot++ on Run 2047
  • 331442 07/11/10 09:13 boyce spot++ on Run 2046
  • 331439 07/11/10 08:29 boyce spot++ of 2045
  • 331435 07/11/10 07:59 mahbub target screen
  • 331434 07/11/10 07:56 mahbub trigger scalers run 2044
  • 331433 07/11/10 07:55 brads Shift Summary (OWL)
  • 331432 07/11/10 07:44 mahbub spot++ run 2044
  • 331431 07/11/10 07:41 brads Bit of a beam excursion at 7:34am
  • 331428 07/11/10 06:51 mahbub spot++ run 2043
  • 331425 07/11/10 06:23 mahbub spot++ run 2042
  • 331424 07/11/10 06:21 mahbub trigger scalers
  • 331423 07/11/10 06:20 brads Hall A Tools
  • 331420 07/11/10 05:25 mahbub spot++ run 2041
  • 331416 07/11/10 04:43 mahbub spot++ run 2040
  • 331413 07/11/10 04:21 brads Updated epicsRunEnd
  • 331412 07/11/10 04:01 mahbub spot++ run 2039
  • 331409 07/11/10 03:49 mahbub target strip chart
  • 331408 07/11/10 03:48 mahbub target screen run 2038
  • 331407 07/11/10 03:35 brads Strange end of run halog entry (331396)
  • 331406 07/11/10 03:30 mahbub trigger scalers run 2038
  • 331398 07/11/10 03:15 brads BPMs and corrector temp snapshot
  • 331395 07/11/10 03:10 brads Beam back, resuming production running (Tools snapshot)
  • 331394 07/11/10 03:10 mahbub spott+ check after beam came back
  • 331392 07/11/10 02:41 N Toro another script to count T6
  • 331391 07/11/10 02:02 brads Hall A Tools snapshot
  • 331390 07/11/10 01:49 brads Septum magnet turned back on
  • 331389 07/11/10 01:42 N Toro right Q1 field after trips
  • 331388 07/11/10 01:30 brads RQ1 back up, Hall C completing their invasive test until 3am
  • 331387 07/11/10 01:20 fansler Right Q1
  • 331386 07/11/10 00:38 brads Tech on-site, investigating Q1 problem
  • 331384 07/11/10 00:14 Gomez Shift summary
  • Saturday

  • 331383 07/10/10 23:32 gomez Hall A in controlled access
  • 331382 07/10/10 23:29 gomez 2nd try - Hall A Tools @23:29 (no beam, RF separator work)
  • 331381 07/10/10 23:23 gomez Hall A Tools @23:22 (no beam, RF separator work)
  • 331380 07/10/10 23:11 Kalyan RF seperator problem
  • 331379 07/10/10 23:11 gomez Q1 right fault#2
  • 331377 07/10/10 22:20 siqbal Spot++ for run #2036
  • 331376 07/10/10 22:18 siqbal Scalars for run #2036
  • 331374 07/10/10 21:38 gomez Q1 right fault
  • 331372 07/10/10 20:46 siqbal Q1 right crashes (run #2034)
  • 331370 07/10/10 20:22 Ole Hansen Re: late end of run logs
  • 331369 07/10/10 20:20 Ole Hansen Re: late end of run logs
  • 331366 07/10/10 19:52 gomez Halog - Late End of Run entries
  • 331365 07/10/10 19:42 siqbal Scalars for run # 2033
  • 331351 07/10/10 17:48 Jack (Techs) RQ1 current control
  • 331349 07/10/10 17:08 siqbal Scalars for run # 2029
  • 331348 07/10/10 16:44 ejensen Shift in coincidence peak
  • 331347 07/10/10 16:36 siqbal Spot++ for run # 2029
  • 331342 07/10/10 16:13 Jixie Zhang Day shift summary
  • 331340 07/10/10 15:59 J. Zhang Increase the field of right Q1 to 0.01416
  • 331336 07/10/10 15:48 K. Bartlett Scaler Screenshot
  • 331335 07/10/10 15:32 J. Zhang Go for 150 micro A beam current (late entry)
  • 331331 07/10/10 15:11 K. Bartlett CODA crashed and rebooted
  • 331330 07/10/10 15:06 J. Zhang Right Q1 was not recover to the previous after fixed
  • 331325 07/10/10 14:44 N Toro turn OFF right Q1 regulation
  • 331323 07/10/10 14:32 K. Bartlett Calibration Runs 2016, 2019
  • 331318 07/10/10 14:19 J. Zhang Q1 tripped and reset
  • 331311 07/10/10 14:10 Kalyan Right arm Q1 field changed.
  • 331310 07/10/10 14:02 Ole Hansen Request one calibration run
  • 331307 07/10/10 13:53 ejensen t1 and t6 coincident peaks for run #2014
  • 331304 07/10/10 13:34 J. Zhang Change A and B laser mixture
  • 331297 07/10/10 13:12 J. Zhang Scripts to track T6 counts
  • 331296 07/10/10 13:09 K. Bartlett Spot++ figures
  • 331292 07/10/10 13:00 K. Bartlett Switched Target to Tantalum
  • 331291 07/10/10 12:53 K. Bartlett Switch Target to Empty
  • 331290 07/10/10 12:32 Jack (Techs) Right Q1 control board replaced
  • 331288 07/10/10 12:03 J. Zhang Septum on
  • 331287 07/10/10 12:02 J. Zhang Go back to beam permission
  • 331276 07/10/10 11:15 K. Bartlett Time Coincidence Distribution for Runs 2004, 2005, 2007
  • 331275 07/10/10 10:08 K. Bartlett Target Screenshots
  • 331274 07/10/10 08:53 J. Zhang Hall is under controlled access
  • 331273 07/10/10 08:49 J. Zhang Left Q1 off
  • 331272 07/10/10 08:15 jmammei Owl Shift Summary
  • 331270 07/10/10 07:38 jmammei attempts to fix RQ1
  • 331269 07/10/10 07:20 jmammei Jack coming in
  • 331268 07/10/10 07:04 jmammei Frequent RQ1 techs
  • 331266 07/10/10 06:43 S.Abrahamyan onlineGUI
  • 331262 07/10/10 06:30 jmammei beam status
  • 331261 07/10/10 06:17 N Toro [techs] Right Q1 tripping frequently, usually requires manual reset
  • 331260 07/10/10 06:04 S.Abrahamyan noisy vdc wires
  • 331259 07/10/10 05:55 jmammei RQ1 trips AGAIN!!!
  • 331256 07/10/10 05:28 jmammei septum on
  • 331254 07/10/10 05:05 jmammei RQ1 trips again
  • 331253 07/10/10 04:59 jmammei dump viewer
  • 331240 07/10/10 03:47 jmammei beam rather stable
  • 331228 07/10/10 03:17 ejensen Noisy left arm vdc wires
  • 331225 07/10/10 02:59 Ole Hansen Swing shift summary (late entry)
  • 331224 07/10/10 02:55 jmammei beam positions
  • 331219 07/10/10 02:26 jmammei beam at 98uA
  • 331218 07/10/10 02:23 jmammei spot++ run 1992
  • 331217 07/10/10 02:18 jmammei beam current 98uA
  • 331214 07/10/10 02:15 jmammei beam back
  • 331212 07/10/10 02:06 Ole Hansen Re: evtypebits
  • 331211 07/10/10 02:04 jmammei beam status
  • 331210 07/10/10 01:57 jmammei paged Jack - not working?
  • 331209 07/10/10 01:56 Ole Hansen Re: deadtime
  • 331208 07/10/10 01:53 ejensen Event type cut explained
  • 331207 07/10/10 01:51 jmammei reset RQ1
  • 331206 07/10/10 01:34 jmammei tech on call not responding
  • 331205 07/10/10 01:26 jmammei RQ1 trips, again.
  • 331204 07/10/10 01:17 ejensen t1 for T6 events
  • 331202 07/10/10 01:10 jmammei crew chief ready to go to 100uA
  • 331201 07/10/10 00:58 ejensen t1 plots for run #1990
  • 331200 07/10/10 00:55 bogdanw look the CODA rate
  • 331198 07/10/10 00:37 jmammei hall access to reset RQ1
  • 331196 07/10/10 00:05 Ole Hansen Right Q1 trip AGAIN
  • 331194 07/10/10 00:01 ejensen Spot++ run #1990
  • Friday

  • 331193 07/09/10 23:55 rouven Improved formatting of Optics Run Summary
  • 331191 07/09/10 23:26 Ole Hansen Ta target, asking for beam
  • 331190 07/09/10 23:26 Ole Hansen Right Q1 reset
  • 331188 07/09/10 23:19 Kalyan Beam current strip chart
  • 331187 07/09/10 23:05 Ole Hansen Right Q1 tripped again
  • 331186 07/09/10 22:40 Ole Hansen High current beam tests underway
  • 331185 07/09/10 22:33 Ole Hansen Carbon bullseye checks
  • 331184 07/09/10 22:11 siqbal Spot++ for run # 1989
  • 331182 07/09/10 22:08 siqbal Pictures for run # 1988
  • 331180 07/09/10 22:04 siqbal Spot++ for run 1988
  • 331175 07/09/10 21:31 ejensen Wierd timing peaks from earlier
  • 331174 07/09/10 21:15 Kalyan Right arm Q1 reset -- procedure to rest
  • 331164 07/09/10 20:06 siqbal Beam turned off
  • 331160 07/09/10 20:02 Ole Hansen R-arm Q1 trip
  • 331159 07/09/10 19:57 Kalyan Beam on Beo viewer
  • 331158 07/09/10 19:28 siqbal Spot++ for run 1980
  • 331146 07/09/10 18:46 Ole Hansen Tune beam in hall
  • 331144 07/09/10 18:45 R. Subedi Day shift summary (late entry)
  • 331121 07/09/10 16:35 Jack (Techs) Right Q3 UNE10 module replaced
  • 331118 07/09/10 15:01 fansler RQ1 lead flow
  • 331117 07/09/10 14:58 R. Subedi Q2,3 of both RHRS cycled
  • 331116 07/09/10 14:50 hsmith Right Dipole (current ramp down issue)
  • 331115 07/09/10 14:47 J. Butler Septum Magnet
  • 331113 07/09/10 14:38 hsmith Right Dipole Polarity change
  • 331109 07/09/10 14:13 Ole Hansen Re: evtypebits
  • 331103 07/09/10 13:40 bogdanw T3/T6
  • 331100 07/09/10 13:35 bogdanw Second PM
  • 331098 07/09/10 13:32 Kalyan Sieve is out
  • 331096 07/09/10 12:09 Bogdan BD magnet
  • 331095 07/09/10 11:58 ejensen Event type cuts
  • 331094 07/09/10 11:57 R. Subedi Hall in controlled access in parallel to pass change, Septum off
  • 331093 07/09/10 11:54 Rouven Essig Summary of Optics Runs
  • 331084 07/09/10 11:25 K. Bartlett Switched to Tantalum Target
  • 331068 07/09/10 10:58 R. Subedi Right Q1 tripped, coming back
  • 331067 07/09/10 10:46 K. Bartlett Target Screenshots
  • 331062 07/09/10 10:19 R. Subedi Right Q1 tripped and brought back
  • 331060 07/09/10 10:15 K. Bartlett Switched to Optics Target
  • 331050 07/09/10 09:45 N Toro dipole is stable at 1.0454
  • 331049 07/09/10 09:40 adaq partial shift summary (July 8 swing)
  • 331048 07/09/10 09:06 schuster moving target to empty
  • 331047 07/09/10 09:05 N Toro set PREX p0 to 1.0454; will post screenshot when dipole is stable
  • 331038 07/09/10 08:42 K. Bartlett Switch to Tantalum
  • 331035 07/09/10 08:27 N Toro owl shift summary (abridged)
  • 331033 07/09/10 08:24 N Toro right q1 down and up again
  • 331024 07/09/10 07:59 N Toro Hall A Tools and magnet stripchart at beginning of run 1934 -- Cycling left quads to 1.16 (right quads fixed since last cycling)
  • 331019 07/09/10 07:52 Neil G-W Runs 1931 and 1932 Junk
  • 331015 07/09/10 07:38 Neil G-W Run 1929 junk
  • 331007 07/09/10 07:10 Neil G-W Correction: Run 1926 Left is 1.1150GeV/c
  • 331004 07/09/10 07:05 N Toro *only* Left HRS at 1.115 GeV/c
  • 331003 07/09/10 06:52 N Toro magnet stripchart: Spectrometer and septum setpoint is 1.160 GeV/c
  • 331002 07/09/10 05:05 Neil G-W Moved to Ta target, beam centered at dump
  • 331001 07/09/10 05:04 N Toro right HRS Q1 down, requesting escorted access
  • 331000 07/09/10 04:49 N Toro moving target to empty for dump viewer test
  • 330999 07/09/10 04:43 Jin Huang Optics Check: dp = -4% Optics Target with Sieve In
  • 330990 07/09/10 04:20 Jin Huang Optics Check: dp=-4% Ta target with sieve in
  • 330987 07/09/10 04:15 N Toro moved target to optics
  • 330978 07/09/10 03:52 V. Sulkosky Tools delta = -4%
  • 330969 07/09/10 03:09 Jin Huang Optics Check: dp=-2% Ta target with sieve in
  • 330957 07/09/10 02:33 N Toro moved to Ta target
  • 330955 07/09/10 02:24 Jin Huang Optics Check: dp=0 optics target with sieve in
  • 330947 07/09/10 02:10 adaq spot++ run 1907
  • 330945 07/09/10 02:00 Neil G-W TO screenshot
  • 330944 07/09/10 01:59 V. Sulkosky Hall A Tools After Q1 Reset
  • 330943 07/09/10 01:58 S.Abrahamyan RHRS VDC problems
  • 330942 07/09/10 01:44 V. Sulkosky Hall Tools at delta = -2%
  • 330941 07/09/10 00:56 Jin Huang Optics Check: dp=0 optics target with sieve in
  • 330939 07/09/10 00:52 N Toro hall A tools snapshot for 0% optics runs
  • Thursday

  • 330924 07/08/10 23:57 Jin Huang Optics Check: Sieve Became Clear with Raster Off
  • 330919 07/08/10 23:44 Jin Huang Sieve with Raster on both arm
  • 330916 07/08/10 23:21 schuster sieves are IN on both arms
  • 330914 07/08/10 22:58 N Toro Moving the central position for optics and reducing raster on optics foil target (BPM (3,-1))
  • 330913 07/08/10 22:56 schuster target change
  • 330902 07/08/10 22:18 Jin Huang What's hit in run 1888
  • 330901 07/08/10 22:17 J. Beacham Carbon optics spot++ screenshots
  • 330891 07/08/10 21:51 Ole Hansen Re: odd dp spectra - evtype cut
  • 330888 07/08/10 21:42 Jin Huang Optics Check: Ta run with raster off
  • 330886 07/08/10 21:34 Ole Hansen Re: R-arm VDC v1 mapping problem?
  • 330882 07/08/10 21:27 V. Sulkosky Re: odd differences between left and right dp spectra
  • 330876 07/08/10 21:13 J. Beacham Moved to carbon optics target
  • 330873 07/08/10 20:58 N Toro above plots were for run #1879
  • 330872 07/08/10 20:53 N Toro odd differences between left and right dp spectra
  • 330867 07/08/10 20:25 J. Beacham Note: Runs 1875 through 1877 have the wrong energies/momenta listed
  • 330865 07/08/10 20:21 N Toro moving target to Tantalum
  • 330864 07/08/10 20:20 J. Beacham Carbon hole centered
  • 330856 07/08/10 19:58 J. Beacham Moved to carbon hole target
  • 330855 07/08/10 19:46 J. Beacham Moved to BeO target
  • 330847 07/08/10 19:14 Kalyan Change to one pass beam
  • 330845 07/08/10 18:25 J. Beacham Moved to empty target
  • 330842 07/08/10 18:16 J. Beacham Carbon hole identified
  • 330834 07/08/10 17:41 J. Beacham Moved to carbon hole target
  • 330833 07/08/10 17:32 J. Beacham Moved to BeO target
  • 330832 07/08/10 17:14 J. Beacham Moved to empty target
  • 330831 07/08/10 17:04 ejensen Day shift summary
  • 330829 07/08/10 16:22 N Toro R-HRS dipole stripchart
  • 330828 07/08/10 16:17 Ole Hansen TDC cuts for evtypebits
  • 330827 07/08/10 16:08 Ole Hansen Re; D.evttypebits examples
  • 330826 07/08/10 16:03 Ole Hansen Re: D.evtypebits examples
  • 330825 07/08/10 15:45 Ole Hansen Re: fEvtHdr.fEvtType (readable)
  • 330824 07/08/10 15:45 Ole Hansen Re: fEvtHdr.fEvtType
  • 330823 07/08/10 15:42 hsmith Right Dipole (current ramp down issue)
  • 330822 07/08/10 15:31 R. Michaels examples of D.evtypebit logic
  • 330821 07/08/10 15:13 hsmith Dipole polarity changed
  • 330820 07/08/10 15:12 R. Michaels re: fEvtHdr.fEvtType vs D.evtypebits
  • 330819 07/08/10 15:10 J. Butler Check List
  • 330817 07/08/10 14:42 ejensen Magnets cycled
  • 330816 07/08/10 14:39 bogdanw lower rate in T5 in low momentum side
  • 330815 07/08/10 14:18 Ole Hansen All VDC spectra posted are for run 1852
  • 330814 07/08/10 14:13 Ole Hansen Re: hamoller reboot partition check
  • 330812 07/08/10 13:57 Ole Hansen R-arm data in trigger 5?!?
  • 330810 07/08/10 13:50 Ole Hansen R-arm VDC wiremaps - clean but bumpy
  • 330809 07/08/10 13:46 Ole Hansen L-arm VDC timing shift with event type
  • 330807 07/08/10 13:39 Ole Hansen L-arm VDC wiremaps - u1 fixed, minor u2 problem
  • 330801 07/08/10 12:18 A. Shahinyan R arm VDC map check
  • 330800 07/08/10 12:15 Neil G-W TO screen capture
  • 330795 07/08/10 11:50 J. Zhang Hamoller rebooted
  • 330779 07/08/10 11:05 folts septum polarity
  • 330778 07/08/10 10:20 ejensen GC #8 delay cable
  • 330773 07/08/10 08:53 Neil G-W Hall is in restricted access
  • 330771 07/08/10 07:58 R. Michaels owl shift summary, part 2
  • 330770 07/08/10 07:45 R. Michaels summary about strange timing concern
  • 330767 07/08/10 06:59 R. Michaels auto abu fixed
  • 330765 07/08/10 05:56 vmaxwell re: L-Arm VDC u2 wire
  • 330762 07/08/10 05:29 vmaxwell TO screenshots
  • 330761 07/08/10 04:41 R. Michaels strange timing affects ADC gate
  • 330758 07/08/10 03:59 Markowitz Owl shift summary, part 1
  • 330757 07/08/10 03:40 Markowitz spot run 1852
  • 330748 07/08/10 03:17 R. Michaels re: strange timing
  • 330747 07/08/10 03:16 vmaxwell L-Arm VDC u2 wire
  • 330745 07/08/10 03:00 Markowitz Trigger rates and data taking rates.
  • 330744 07/08/10 03:00 Markowitz Trigger rates and data taking rates.
  • 330740 07/08/10 02:46 vmaxwell strange timing for run #1850
  • 330738 07/08/10 01:25 R. Michaels Empty histograms
  • 330737 07/08/10 01:00 R. Michaels epicsRunStart was slow, now faster
  • 330736 07/08/10 00:52 vmaxwell 1849 is production run
  • 330733 07/08/10 00:39 Markowitz spot++ run 1849
  • 330730 07/08/10 00:22 R. Michaels moving to Ta tgt
  • 330729 07/08/10 00:20 Markowitz spot++
  • 330726 07/08/10 00:06 R. Michaels target to BeO
  • 330724 07/08/10 00:00 LeRose Shift Summary
  • Wednesday

  • 330723 07/07/10 23:53 LeRose Tune Beam
  • 330722 07/07/10 23:42 J. Beacham Moved to empty target
  • 330701 07/07/10 22:10 siqbal Cosmic runs
  • 330700 07/07/10 22:05 J. Beacham TO screenshots
  • 330699 07/07/10 22:04 LeRose Beam in >= 1 hour
  • 330695 07/07/10 21:42 LeRose Waiting for Beam
  • 330694 07/07/10 20:02 LeRose Spectrometers Restored
  • 330693 07/07/10 20:00 LeRose Power Reduction Ended
  • 330685 07/07/10 18:58 ejensen GC #8 delay cable
  • 330682 07/07/10 14:15 Ole Hansen Day Shift Summary (reformatted)
  • 330681 07/07/10 14:01 David Anez TO Screenshots
  • 330680 07/07/10 13:59 Ole Hansen Day Shift Summary
  • 330679 07/07/10 13:18 Ole Hansen Beam off for power reduction period
  • 330677 07/07/10 12:50 Ole Hansen Bad VDC amplifier card L-arm u1-plane
  • 330676 07/07/10 12:35 David Anez Spot++, Run 1832, Tantalum Target
  • 330674 07/07/10 12:29 David Anez Spot++, Run 1798, BeO Target
  • 330673 07/07/10 12:26 Ole Hansen Resuming production
  • 330672 07/07/10 12:26 David Anez Spot++, Run 1799, BeO Target
  • 330671 07/07/10 12:23 David Anez Spot++, Run 1800, Carbon Hole Target
  • 330670 07/07/10 12:21 David Anez Spot++, Run 1801, Carbon Hole Target
  • 330669 07/07/10 12:20 David Anez Spot++, Run 1802, Carbon Hole Target
  • 330668 07/07/10 12:19 David Anez Spot++, Run 1803, Carbon Hole Target
  • 330667 07/07/10 12:17 David Anez Spot++, Run 1804, Carbon Hole Target
  • 330666 07/07/10 12:16 Ole Hansen Restarted adaql7 printer queue
  • 330665 07/07/10 12:15 David Anez Spot++, Run 1805, Carbon Hole Target
  • 330664 07/07/10 12:15 David Anez Moved to Tantalum Target
  • 330663 07/07/10 12:11 David Anez Spot++, Run 1806, Optics Target
  • 330662 07/07/10 12:10 David Anez Spot++, Run 1807, Optics Target
  • 330661 07/07/10 12:08 David Anez Spot++, Run 1808, Tantalum Target
  • 330660 07/07/10 12:06 David Anez Spot++, Run 1809, Tantalum Target
  • 330659 07/07/10 11:54 David Anez Target moved to Empty
  • 330658 07/07/10 11:54 Ole Hansen Requesting dump check
  • 330656 07/07/10 11:47 David Anez Spot++, Run 1810, Tantalum Target
  • 330655 07/07/10 11:45 David Anez Spot++, Run 1811, Tantalum Target
  • 330654 07/07/10 11:43 David Anez Spot++, Run 1812, Tantalum Target
  • 330653 07/07/10 11:42 David Anez Spot++, Run 1813, Tantalum Target
  • 330652 07/07/10 11:41 David Anez Spot++, Run 1814, Tantalum Target
  • 330651 07/07/10 11:34 David Anez Spot++, Run 1815, Tantalum Target
  • 330650 07/07/10 11:33 David Anez Spot++, Run 1816, Tantalum Target
  • 330649 07/07/10 11:31 David Anez Spot++, Run 1817, Tantalum Target
  • 330648 07/07/10 11:30 David Anez Spot++, Run 1818, Tantalum Target
  • 330647 07/07/10 11:28 David Anez Spot++, Run 1819, Tantalum Target
  • 330646 07/07/10 11:27 David Anez Spot++, Run 1820, Tantalum Target
  • 330645 07/07/10 11:25 David Anez Spot++, Run 1821, Tantalum Target
  • 330644 07/07/10 11:22 David Anez Spot++, Run 1822, Tantalum Target
  • 330635 07/07/10 11:17 David Anez Spot++, Run 1824, Tantalum Target (correct)
  • 330634 07/07/10 11:15 bono missing shower histogram
  • 330633 07/07/10 11:00 David Anez Spot++, Run 1826, Tantalum Target
  • 330632 07/07/10 10:58 David Anez Spot++, Run 1824, Tantalum Target
  • 330631 07/07/10 10:54 David Anez Spot++, Run 1831, Tantalum Target
  • 330629 07/07/10 10:46 Ole Hansen Disabling spurious HRS detector alarms
  • 330628 07/07/10 10:32 David Anez Spot++, Run 1830, Tantalum Target
  • 330627 07/07/10 10:29 David Anez Spot++, Run 1829, Tantalum Target
  • 330625 07/07/10 10:15 David Anez TO Screenshots
  • 330621 07/07/10 09:17 bono this is for 1828 not 1829
  • 330620 07/07/10 09:09 bono spot_L run=1827 tantalum target
  • 330617 07/07/10 09:04 bono spot_L run=1828 tantalum target
  • 330616 07/07/10 08:51 bono spot_L run=1826 tantalum target
  • 330614 07/07/10 08:08 Markowitz Owl shift summary
  • 330611 07/07/10 07:46 ejensen Run start/end log entries
  • 330610 07/07/10 07:42 vmaxwell TO screenshots
  • 330609 07/07/10 07:06 ejensen Run start/end log entries
  • 330607 07/07/10 06:54 ejensen Updated left arm s2m pedestals
  • 330605 07/07/10 06:09 S.Abrahamyan ShowerPreShower calibration
  • 330604 07/07/10 06:01 S.Abrahamyan Shower/PreShower calibration
  • 330602 07/07/10 05:48 ejensen Pion rejection plots
  • 330600 07/07/10 05:37 ejensen Right HRS GC pion rejection efficiency
  • 330599 07/07/10 05:11 adaq spot run 1825
  • 330597 07/07/10 04:58 ejensen t1 plot run #1824
  • 330593 07/07/10 03:41 Markowitz spot run 1821
  • 330592 07/07/10 03:39 Markowitz spot run 1821 -- scraping?
  • 330590 07/07/10 03:33 ejensen Spot++ new raster size 8x6
  • 330589 07/07/10 03:31 ejensen Spot++ new raster size 8x6
  • 330588 07/07/10 03:31 ejensen Raster looks odd
  • 330577 07/07/10 03:16 A. Camsonne Rim 1813 - target y distribution
  • 330573 07/07/10 03:00 Jin Huang Optics Check: check for RHRS sieve-in run with T6 trigger
  • 330572 07/07/10 02:57 A. Camsonne Restored coincidence trigger T6
  • 330571 07/07/10 02:52 Jin Huang Optics Check: Locating foils on Ta target
  • 330570 07/07/10 02:46 Markowitz Access to restore T6 trigger
  • 330567 07/07/10 02:38 Jin Huang Optics Check: sieve more clear for L-HRS sieve-in run with T5 trigger
  • 330564 07/07/10 02:17 Jin Huang Optics Check: Locating foils on optics target
  • 330562 07/07/10 01:54 N Toro sieve partially in puzzle
  • 330559 07/07/10 01:11 Markowitz Turn septum back on.
  • 330558 07/07/10 01:10 Markowitz Sieve out access
  • 330555 07/07/10 00:37 Markowitz Septum back on.
  • 330554 07/07/10 00:36 Markowitz Call for an access to take sieve half-out
  • 330551 07/07/10 00:12 vmaxwell TO Screenshots
  • 330549 07/07/10 00:01 N Toro optics runs / last hr of swing shift
  • Tuesday

  • 330541 07/06/10 23:23 N Toro on access: sieve in, T6->positron singles
  • 330540 07/06/10 23:14 Kalyan Swing shift summary
  • 330538 07/06/10 22:37 boyce spot++ on Tantalum 1809
  • 330537 07/06/10 22:35 A. Camsonne Optics foils - Run 1806
  • 330533 07/06/10 22:06 J. Beacham Moved to tantalum target
  • 330528 07/06/10 21:44 riordan Round BeO Spot with Centered Carbon Hole
  • 330527 07/06/10 21:40 J. Beacham Moved to carbon optics target
  • 330526 07/06/10 21:30 J. Beacham Moved to BeO target
  • 330525 07/06/10 21:28 boyce Finally found hole in C target
  • 330512 07/06/10 20:56 riordan BPM Positions for BeO
  • 330511 07/06/10 20:54 riordan Spot looks much better!
  • 330510 07/06/10 20:54 J. Beacham Moved to carbon hole target
  • 330509 07/06/10 20:17 riordan Spot On BeO has vertical spread
  • 330508 07/06/10 20:16 A. Camsonne Quad was off design since June 28th
  • 330502 07/06/10 19:32 J. Beacham Moved to BeO target
  • 330493 07/06/10 19:08 riordan RQ1 tripped again - reset from GUI
  • 330492 07/06/10 18:44 J. Beacham Moved to empty target
  • 330489 07/06/10 17:34 S.Abrahamyan Leadglass HV
  • 330484 07/06/10 16:58 J. Beacham TO screenshots
  • 330483 07/06/10 16:19 Jin Huang Optics Check: LHRS reconstruction of Ta foil vz, before and after adding bending magnets
  • 330482 07/06/10 16:08 A. Camsonne Target light put on hacweb8 instead of heater
  • 330480 07/06/10 15:35 S.Abrahamyan Shower/PreShower HV
  • 330479 07/06/10 15:21 hsmith Left Q1
  • 330478 07/06/10 14:02 David Anez TO Screenshots
  • 330477 07/06/10 13:20 A. Shahinyan S2m Left PMT's cabling in HRS-R
  • 330472 07/06/10 12:00 David Anez TO Screenshots
  • 330471 07/06/10 11:58 David Anez Spot++, Run 1790, Tantalum Target
  • 330470 07/06/10 11:57 David Anez Spot++, Run 1789, Tantalum Target
  • 330469 07/06/10 11:55 David Anez Spot++, Run 1788, Tantalum Target
  • 330468 07/06/10 11:54 David Anez Spot++, Run 1787, Tantalum Target
  • 330467 07/06/10 11:53 David Anez Spot++, Run 1786, Lead 1 Target
  • 330466 07/06/10 11:51 David Anez Spot++, Run 1785, Lead 1 Target
  • 330465 07/06/10 11:50 David Anez Spot++, Run 1784, Lead 1 Target
  • 330464 07/06/10 11:49 David Anez Spot++, Run 1783, Tantalum Target
  • 330463 07/06/10 11:47 David Anez Spot++, Run 1780, Tantalum Target
  • 330462 07/06/10 11:46 David Anez Spot++, Run 1779, Tantalum Target
  • 330461 07/06/10 11:44 David Anez Spot++, Run 1778, Tantalum Target
  • 330460 07/06/10 11:43 David Anez Spot++, Run 1777, Tantalum Target
  • 330459 07/06/10 11:40 David Anez Spot++, Run 1776, Tantalum Target
  • 330458 07/06/10 11:39 David Anez Spot++, Run 1775, Tantalum Target
  • 330457 07/06/10 11:38 David Anez Spot++, Run 1774, Tantalum Target
  • 330456 07/06/10 11:37 David Anez Spot++, Run 1773, Tantalum Target
  • 330455 07/06/10 11:36 David Anez Spot++, Run 1772, Tantalum Target
  • 330454 07/06/10 11:34 David Anez Spot++, Run 1771, Tantalum Target
  • 330453 07/06/10 11:33 David Anez Spot++, Run 1770, Tantalum Target
  • 330452 07/06/10 11:32 David Anez Spot++, Run 1769, Tantalum Target
  • 330451 07/06/10 11:30 David Anez Spot++, Run 1768, Tantalum Target
  • 330450 07/06/10 11:29 David Anez Spot++, Run 1767, Thick Carbon Target
  • 330449 07/06/10 11:27 David Anez Spot++, Run 1766, Thick Carbon Target
  • 330448 07/06/10 11:26 David Anez Spot++, Run 1765, Tantalum Target
  • 330447 07/06/10 11:24 David Anez Spot++, Run 1764, Optics Target
  • 330446 07/06/10 11:22 David Anez Spot++, Run 1763, Optics Target
  • 330445 07/06/10 11:21 David Anez Spot++, Run 1762, Optics Target
  • 330444 07/06/10 11:20 David Anez Spot++, Run 1761, Optics Target
  • 330436 07/06/10 11:19 David Anez Spot++, Run 1760, Optics Target
  • 330435 07/06/10 11:18 David Anez Spot++, Run 1759, Optics Target
  • 330434 07/06/10 11:17 David Anez Spot++, Run 1758, Optics Target
  • 330433 07/06/10 11:15 David Anez Spot++, Run 1757, Optics Target
  • 330432 07/06/10 11:14 David Anez Spot++, Run 1756, Optics Target
  • 330431 07/06/10 11:12 David Anez Spot++, Run 1755, Optics Target
  • 330430 07/06/10 11:11 David Anez Spot++, Run 1754, Optics Target
  • 330429 07/06/10 11:10 David Anez Spot++, Run 1753, Optics Target
  • 330428 07/06/10 11:08 David Anez Spot++, Run 1752, Optics Target
  • 330427 07/06/10 11:03 David Anez Spot++, Run 1751, Lead 1 Target
  • 330426 07/06/10 11:00 David Anez Spot++, Run 1750, Lead 1 Target
  • 330425 07/06/10 10:59 David Anez Spot++, Run 1749, Lead 1 Target
  • 330424 07/06/10 10:58 David Anez Spot++, Run 1748, Optics Target
  • 330423 07/06/10 10:57 David Anez Spot++, Run 1747, Optics Target
  • 330422 07/06/10 10:56 David Anez Spot++, Run 1746, Optics Target
  • 330421 07/06/10 10:55 David Anez Spot++, Run 1745, Optics Target
  • 330420 07/06/10 10:52 David Anez Spot++, Run 1744, Optics Target
  • 330419 07/06/10 10:50 David Anez Spot++, Run 1743, Optics Target
  • 330418 07/06/10 10:49 David Anez Spot++, Run 1742, Tantalum Target
  • 330417 07/06/10 10:47 David Anez Spot++, Run 1741, Tantalum Target
  • 330416 07/06/10 10:46 David Anez Spot++, Run 1740, Tantalum Target
  • 330415 07/06/10 10:45 David Anez Spot++, Run 1739, Tantalum Target
  • 330414 07/06/10 10:44 David Anez Spot++, Run 1738, Tantalum Target
  • 330413 07/06/10 10:43 David Anez Spot++, Run 1737, Tantalum Target
  • 330412 07/06/10 10:41 David Anez Spot++, Run 1736, Tantalum Target
  • 330411 07/06/10 10:40 David Anez Spot++, Run 1735, Tantalum Target
  • 330410 07/06/10 10:39 David Anez Spot++, Run 1734, Tantalum Target
  • 330409 07/06/10 10:38 David Anez Spot++, Run 1733, Tantalum Target
  • 330408 07/06/10 10:37 David Anez Spot++, Run 1732, Lead 1 Target
  • 330407 07/06/10 10:35 David Anez Spot++, Run 1731, Lead 1 Target
  • 330406 07/06/10 10:34 David Anez Spot++, Run 1730, Lead 1 Target
  • 330405 07/06/10 10:33 LeRose Magnets Look OK
  • 330404 07/06/10 10:32 David Anez Spot++, Run 1729, Lead 1
  • 330403 07/06/10 10:27 David Anez Spot++, Run 1728, Lead 1 Target
  • 330402 07/06/10 10:23 David Anez Incorrect Spot++ in 330400 for Run 1728
  • 330401 07/06/10 10:10 David Anez TO Screenshots
  • 330400 07/06/10 10:07 David Anez 1728 screenshots
  • 330399 07/06/10 07:57 lselvy Access
  • 330398 07/06/10 07:50 lselvy Screenshots
  • 330396 07/06/10 06:02 lselvy Screenshots
  • 330395 07/06/10 06:02 Markowitz Owl shift summary
  • 330393 07/06/10 05:53 vmaxwell Timing Peaks
  • 330392 07/06/10 05:00 Markowitz spot++
  • 330389 07/06/10 04:06 lselvy Screenshots
  • 330378 07/06/10 02:17 Markowitz Access to reset Q1 done; septum back on
  • 330377 07/06/10 02:10 Markowitz RHRS Q1 trips off
  • 330374 07/06/10 01:50 ejensen Tantalum peak (no cuts)
  • 330373 07/06/10 01:49 N Toro Timing peak in Pb run 1786
  • 330372 07/06/10 01:44 Markowitz spot++, run 1787 on Ta
  • 330371 07/06/10 01:42 Markowitz BCM monitors are crazily uncalibrated
  • 330370 07/06/10 01:38 Markowitz changing to Ta target
  • 330369 07/06/10 01:35 lselvy Screenshots
  • 330367 07/06/10 01:35 lselvy Target Changed
  • 330366 07/06/10 01:05 Markowitz spot++, run 1786
  • 330365 07/06/10 00:58 Markowitz LHRS Q1 oscillating
  • 330363 07/06/10 00:44 Markowitz raster to 4x4 and beam spot y =1 mm
  • 330358 07/06/10 00:32 ejensen Tantalum peak
  • 330357 07/06/10 00:26 N Toro target is Pb/C#1; septum is back on
  • 330355 07/06/10 00:01 N Toro July 5 Swing shift summary
  • Monday

  • 330354 07/05/10 23:57 N Toro q1 needs reset; requesting controlled access
  • 330352 07/05/10 23:45 A. Camsonne Problem with replay - work disk full - removing old PREX root files
  • 330348 07/05/10 23:06 Jin Huang Optics Check: sieve FOUND for L-HRS sieve-in run with T5 trigger
  • 330345 07/05/10 23:04 N Toro decreased beam current to 1uA, raster to 6x6 MCC
  • 330344 07/05/10 23:00 adaq spot++ run 1778 on Ta
  • 330343 07/05/10 22:59 Ole Hansen Re: R-arm VDC performance/issues
  • 330340 07/05/10 22:39 Jin Huang Optics Check: more check for sieve-in run with T6 trigger
  • 330339 07/05/10 22:33 A. Camsonne VDC Argon cylinder swapped
  • 330335 07/05/10 22:24 N Toro raster is 2x2 MCC
  • 330331 07/05/10 22:06 N Toro sieve is OUT, T6 two-arm coincidence restored
  • 330328 07/05/10 21:55 Jin Huang Optics Check: more statistics for sieve-in run with T6 trigger
  • 330327 07/05/10 21:36 adaq septum off for access
  • 330322 07/05/10 21:14 Jin Huang Optics Check: quick look on sieve in run with T6 trigger
  • 330320 07/05/10 21:09 ejensen t1 for run #1767
  • 330316 07/05/10 20:56 Ole Hansen R-arm VDC performace/issues
  • 330311 07/05/10 20:41 Jin Huang Optics Check: Locating foils on optics target
  • 330310 07/05/10 20:37 N Toro T6 modified temporarily to single-arm positron trigger
  • 330309 07/05/10 20:35 schuster sieve is in
  • 330308 07/05/10 20:32 N Toro septum back on
  • 330307 07/05/10 20:21 N Toro moving to tantalum target for sieve-in optics run
  • 330306 07/05/10 20:14 N Toro septum off for escorted access
  • 330294 07/05/10 19:14 ZAFAR AHMED Target is thick carbon
  • 330291 07/05/10 19:10 N Toro t1 timing (25 uA on Ta)
  • 330290 07/05/10 19:05 ZAFAR AHMED Target is Tantalum
  • 330275 07/05/10 18:15 ZAFAR AHMED Spot++ runs
  • 330266 07/05/10 17:49 ZAFAR AHMED Spot++
  • 330262 07/05/10 17:38 schuster move target to optics
  • 330259 07/05/10 17:29 Jin Huang RHRS Optics Run Checks: Optics Target VZ
  • 330258 07/05/10 17:17 Ole Hansen R-arm VDC database TDC offsets updated
  • 330256 07/05/10 17:04 Ole Hansen Re; Optics check runs
  • 330254 07/05/10 16:33 N Toro coincidence run on Pb #1
  • 330251 07/05/10 16:15 Kalyan Day Shift Summary
  • 330247 07/05/10 16:02 David Anez Target moved to Lead 1
  • 330246 07/05/10 16:01 David Anez Optics Check runs
  • 330237 07/05/10 15:28 Ole Hansen R-arm VDC database updated
  • 330232 07/05/10 15:16 Ole Hansen Low VDC hit efficiency run 1735
  • 330231 07/05/10 15:15 David Anez Target moved to optics
  • 330227 07/05/10 14:48 David Anez Raster ON - Runs 1736-1740
  • 330225 07/05/10 14:33 Jin Huang RHRS Optics Run Checks
  • 330223 07/05/10 14:21 David Anez TO Screenshots
  • 330209 07/05/10 13:07 ejensen s2m cable puzzle
  • 330206 07/05/10 12:39 David Anez Target moved to Ta
  • 330203 07/05/10 12:36 David Anez Runs 1730,1731 information
  • 330202 07/05/10 12:31 ejensen Correction: ADC gate
  • 330199 07/05/10 12:09 David Anez Screenshots
  • 330198 07/05/10 11:50 riordan RQ1 Reset
  • 330197 07/05/10 11:24 Kalyan Right arm Q1 tripped
  • 330185 07/05/10 11:07 ejensen Timing Gate
  • 330184 07/05/10 11:01 ejensen Right Arm Delay Cables for cherenkov and s2m
  • 330183 07/05/10 10:30 Kalyan Hall in beam permit but there is still no beam
  • 330182 07/05/10 10:20 David Anez Screenshots
  • 330181 07/05/10 09:41 Kalyan Hall in controlled access
  • 330179 07/05/10 08:39 Kalyan The beam is still down
  • 330178 07/05/10 08:21 David Anez Screenshots
  • 330176 07/05/10 08:05 Markowitz Owl shift summary
  • 330175 07/05/10 07:31 A. Camsonne Positrons in coincidence Run 1723
  • 330174 07/05/10 07:08 A. Camsonne Coincidence peak with 40 ns coincidence gate - Run 1723
  • 330172 07/05/10 06:40 Markowitz Optics runs
  • 330164 07/05/10 06:07 Markowitz Sieve out
  • 330163 07/05/10 06:06 Markowitz Giving up on optics at 2 pass
  • 330161 07/05/10 05:55 S.Abrahamyan Sower/PreShower HV
  • 330159 07/05/10 05:49 lselvy Screenshots
  • 330155 07/05/10 05:17 A. Camsonne Left Arm HV
  • 330147 07/05/10 04:42 Markowitz Q1 tripped
  • 330144 07/05/10 04:36 B. Wojtsekho Gate width for T4 and T6 reduced
  • 330143 07/05/10 04:23 A. Camsonne L1A scaler plugged back
  • 330141 07/05/10 04:14 lselvy Target Change
  • 330138 07/05/10 04:07 lselvy Screenshots
  • 330136 07/05/10 03:55 J. Beacham Controlled access
  • 330135 07/05/10 03:41 lselvy Target Change
  • 330127 07/05/10 02:48 lselvy Target Change
  • 330126 07/05/10 02:47 A. Camsonne [Late Entry] Q1 power supply cycled
  • 330125 07/05/10 02:39 lselvy Target Change
  • 330124 07/05/10 02:28 lselvy Target Change
  • 330122 07/05/10 02:16 lselvy Screenshots
  • 330121 07/05/10 02:16 A. Camsonne Odd structure in T1 timing spectrum
  • 330117 07/05/10 00:52 J. Beacham Controlled access
  • 330116 07/05/10 00:29 lselvy Target Change
  • 330115 07/05/10 00:27 A. Camsonne Rough estimation signal over background
  • 330114 07/05/10 00:18 lselvy Target Change
  • 330113 07/05/10 00:17 bogdanw trigger work
  • 330112 07/05/10 00:14 A. Camsonne T1 timing spectrum for run 1702 ( delayed 100 ns ) and 1645
  • Sunday

  • 330109 07/04/10 23:58 lselvy Vacuum Screenshot
  • 330105 07/04/10 23:54 N Toro Shift summary (swing)
  • 330103 07/04/10 23:54 N Toro Q1 back up
  • 330101 07/04/10 23:51 N Toro turned off septum for controlled access
  • 330100 07/04/10 23:23 adaq HRS-R S0 HV
  • 330090 07/04/10 23:03 Ole Hansen Re: VDC database
  • 330086 07/04/10 22:59 A. Camsonne Timing spectra for run 1702 low statistics
  • 330081 07/04/10 22:39 N Toro septum on
  • 330080 07/04/10 22:31 A. Camsonne 100 ns delay added to Left S2m signal by Bogdan
  • 330079 07/04/10 22:25 A. Camsonne VDC database
  • 330078 07/04/10 22:21 A. Camsonne Added shower histograms to two arm replay
  • 330075 07/04/10 22:03 A. Camsonne VDC database issue
  • 330067 07/04/10 21:47 adaq scalers at 1uA
  • 330061 07/04/10 21:29 riordan BeO Scan Results
  • 330060 07/04/10 21:28 David Anez Moved to Lead 1 Target
  • 330055 07/04/10 20:12 riordan Septum ON
  • 330054 07/04/10 20:07 David Anez move target beam
  • 330039 07/04/10 17:58 riordan CHL Update
  • 330031 07/04/10 16:08 Kalyan Day shift summary
  • 330028 07/04/10 15:51 Kalyan Right arm sieve in run : 1671
  • 330025 07/04/10 14:57 ZAFAR AHMED Target strip chart
  • 330024 07/04/10 14:56 ZAFAR AHMED Target GUI
  • 330023 07/04/10 14:47 riordan VDC Time Offsets Updated
  • 330021 07/04/10 14:25 riordan Sieve is OUT for left and right
  • 330020 07/04/10 14:23 Kalyan Hall in controlled access
  • 330019 07/04/10 14:15 Kalyan CHL tripped
  • 330010 07/04/10 13:43 ZAFAR AHMED Target is empty target
  • 330007 07/04/10 13:29 R. Michaels Trigger 6 coincidence time spectrum
  • 330006 07/04/10 13:27 ZAFAR AHMED Target is BeO
  • 330005 07/04/10 13:17 A. Camsonne Threshold readback not working
  • 330003 07/04/10 13:11 Ole Hansen VDC threshold setpoints
  • 330002 07/04/10 13:11 ZAFAR AHMED Sieve runs
  • 329999 07/04/10 13:07 ZAFAR AHMED Spot++ at (0,0)
  • 329984 07/04/10 12:19 Kalyan Back to beam permit
  • 329983 07/04/10 11:44 riordan Sieve is IN for Left and Right
  • 329982 07/04/10 11:42 ejensen Threshold arrangement for VDC
  • 329981 07/04/10 11:42 A. Camsonne Gas cerenkov threshold
  • 329980 07/04/10 11:37 Kalyan MCC finished ion chamber calibration
  • 329974 07/04/10 11:25 Kalyan MCC is calibrating the ion chambers
  • 329955 07/04/10 10:44 Kalyan R-HRS Q1 tripped
  • 329954 07/04/10 10:42 Ole Hansen VDC threshold
  • 329953 07/04/10 10:32 Kalyan Hall in beam permit
  • 329952 07/04/10 10:30 ejensen Right arm s2m HV adjustment
  • 329943 07/04/10 08:59 Kalyan Hall in controlled access
  • 329939 07/04/10 08:15 Markowitz Owl shift summary: July 4, 2010
  • 329934 07/04/10 07:45 lselvy Sreenshots
  • 329933 07/04/10 07:43 J. Beacham spot++ -- raster may be scraping the edge
  • 329928 07/04/10 07:32 J. Beacham spot++
  • 329926 07/04/10 07:24 Markowitz Decrease raster size
  • 329924 07/04/10 07:16 J. Beacham spot++
  • 329921 07/04/10 06:47 Markowitz Further S0 HV adjustment
  • 329920 07/04/10 06:36 J. Beacham spot++
  • 329918 07/04/10 06:25 Markowitz Beam on dump viewer verified
  • 329917 07/04/10 06:25 lselvy Target Change
  • 329916 07/04/10 06:18 lselvy Target Change
  • 329915 07/04/10 06:16 Markowitz septum back on
  • 329912 07/04/10 06:04 Markowitz So HV to 1400
  • 329906 07/04/10 05:46 vmaxwell HRS right arm HV changes
  • 329905 07/04/10 05:44 S.Abrahamyan S2M timing
  • 329903 07/04/10 05:20 Markowitz septum off for access
  • 329901 07/04/10 04:57 lselvy Screenshots
  • 329888 07/04/10 03:56 J. Beacham spot++
  • 329883 07/04/10 03:44 J. Beacham spot++
  • 329882 07/04/10 03:43 Markowitz tools screenshots
  • 329881 07/04/10 03:41 lselvy Vacuum Screenshot
  • 329880 07/04/10 03:41 Markowitz RHRS Q1 trips off. Software reset works.
  • 329877 07/04/10 03:31 lselvy Target Change
  • 329876 07/04/10 03:29 J. Beacham spot++ empty target
  • 329872 07/04/10 03:19 Markowitz Turning on septum
  • 329869 07/04/10 03:11 lselvy Target
  • 329829 07/04/10 00:12 vmaxwell current runs
  • 329828 07/04/10 00:09 N Toro Shift summary (Swing)
  • Saturday

  • 329807 07/03/10 22:03 N Toro septum off; controlled access
  • 329806 07/03/10 21:57 N Toro spot++ moller quads back up
  • 329799 07/03/10 21:22 A. Camsonne [techs] Q1 back up
  • 329797 07/03/10 20:55 A. Camsonne Right arm detector issues so far - From run 1553
  • 329795 07/03/10 20:51 A. Camsone Odd target Y distribution
  • 329794 07/03/10 20:49 N Toro VDC HV to 3.5 kV
  • 329793 07/03/10 20:39 A. Camsonne [techs] Q1 down Phase/Unbalance failure
  • 329792 07/03/10 20:39 ejensen End of run log entries
  • 329789 07/03/10 20:16 ejensen Right arm shower and preshower pedestals
  • 329788 07/03/10 20:16 N Toro high-current test of permanent magnet temperature is complete
  • 329786 07/03/10 20:14 Schuster Detector checkout
  • 329785 07/03/10 20:05 David Anez move lead
  • 329784 07/03/10 19:54 N Toro end of temperature test, moving to empty
  • 329783 07/03/10 19:54 David Anez move empty target
  • 329782 07/03/10 19:53 A. Camsonne Right arm gas cerenkov change
  • 329777 07/03/10 19:44 R. Michaels comments about CT*f4 in run 1552
  • 329760 07/03/10 19:13 N Toro permanent magnet heating test so far
  • 329751 07/03/10 18:59 ejensen s2m pedestals updated
  • 329745 07/03/10 18:51 N Toro VDC HV is 3kV
  • 329744 07/03/10 18:50 N Toro starting magnet high-current test
  • 329742 07/03/10 18:46 N Toro lowered VDC HV to 3V
  • 329741 07/03/10 18:43 N Toro stripchart of magnet cooling
  • 329740 07/03/10 18:41 N Toro beam on dump at 67.5, beginning high-current permanent magnet temperature test
  • 329739 07/03/10 18:41 David Anez beam move lead target
  • 329738 07/03/10 18:32 N Toro MCC moving beam to Y=+1 at 1H01
  • 329737 07/03/10 18:32 David Anez move target
  • 329736 07/03/10 18:30 A. Camsonne Pulser run for pedestals run 1578
  • 329735 07/03/10 18:29 N Toro magnet heating test
  • 329731 07/03/10 18:25 A. Camsonne Right arm S2m amplitudes seems fine
  • 329728 07/03/10 18:16 A. Camsonne Argon and CO2 cylinders swapped
  • 329727 07/03/10 18:14 ejensen Right arm gas cherenkov 8 and 9
  • 329717 07/03/10 17:52 A. Camsonne VDC current draw at 5uA
  • 329715 07/03/10 17:51 ejensen s0 for right arm
  • 329714 07/03/10 17:49 ejensen Shower 36 on right arm
  • 329712 07/03/10 17:42 riordan spot++ 16x16 raster
  • 329710 07/03/10 17:36 A. Camsonne Trigger rates at 0.7 uA on Pb1
  • 329709 07/03/10 17:34 A. Camsonne Triggers timing with L and R cerenkov cut
  • 329693 07/03/10 17:05 A. Camsonne Ct1f4 and Ct3f4 run 1552
  • 329690 07/03/10 16:43 R. Michaels L1A not plugged into scaler ?
  • 329687 07/03/10 16:36 A. Camsonne Re : Call in
  • 329686 07/03/10 16:34 A. Camsonne Correlation timing T1 and timing T3
  • 329682 07/03/10 16:28 R. Lindgren Shift Summary
  • 329670 07/03/10 15:46 A. Camsonne Peaks in timing spectrum run 1552
  • 329669 07/03/10 15:44 folts call in
  • 329668 07/03/10 15:09 N Toro beam sometimes hits magnet?
  • 329662 07/03/10 14:47 Lindgren Flashing light
  • 329661 07/03/10 14:41 A. Camsonne Confirmed that the flashing is a light
  • 329659 07/03/10 14:22 L:indgren Flashing light
  • 329652 07/03/10 13:33 Lindgren Target change
  • 329651 07/03/10 13:31 V.Owen target moved
  • 329645 07/03/10 12:43 Lindgren Corrector magnet tripping
  • 329640 07/03/10 12:11 R. Michaels T4 coinc plots
  • 329617 07/03/10 11:04 V.Owen target moved
  • 329610 07/03/10 10:41 riordan spot++, Raster 12x8 MCC units
  • 329608 07/03/10 10:35 riordan spot++, 4x4 MCC Raster
  • 329607 07/03/10 10:34 Lindgren Beam in Hall A
  • 329603 07/03/10 10:13 R. Michaels coincidence search, but problem
  • 329602 07/03/10 09:52 R. Lindgren No Beam yet
  • 329597 07/03/10 08:48 A. Camsonne Plot T1 vs T3 for run 1510
  • 329596 07/03/10 08:39 A. Camsonne Timing from S0 run 1510
  • 329595 07/03/10 08:37 Zafar Ahmed/ Incomplete Shift Summary
  • 329594 07/03/10 08:31 A. Camsonne Run 1519 trigger timing spectra.
  • 329592 07/03/10 08:28 A. Camsonne Run 1510 timing trigger 1 vs trigger 3
  • 329591 07/03/10 08:14 A. Camsonne Right arm S0 database modified
  • 329587 07/03/10 08:07 A. Camsonne Run 1519 simulated coincidence run
  • 329581 07/03/10 07:34 I. Pomerantz RHRS s2m HV calibration
  • 329580 07/03/10 07:29 A. Camsonne Trigger latch on Left arm
  • 329579 07/03/10 07:25 A. Camsonne D class database updated in 20100701
  • 329571 07/03/10 06:58 A. Camsonne Right scalers display
  • 329568 07/03/10 06:46 riordan spot++ doesn't look right
  • 329566 07/03/10 05:48 N Toro status of beamline commissioning
  • 329565 07/03/10 05:46 A. Camsonne Page received
  • 329564 07/03/10 05:39 N Toro powering down septum
  • 329563 07/03/10 05:33 N Toro beamline commissioning log
  • 329562 07/03/10 05:32 N Toro going to controlled access
  • 329561 07/03/10 05:26 Kiadtisak Target Moved to Empty
  • 329559 07/03/10 05:23 N Toro spot++ 4x4MCC carbon hole target
  • 329552 07/03/10 05:01 S.Abrahamyan VDC
  • 329546 07/03/10 04:51 Kiadtisak Target moved to Carbon Hole
  • 329545 07/03/10 04:48 adaq both bpms to 0,0
  • 329543 07/03/10 04:09 N Toro septum on 775.6 A
  • 329541 07/03/10 04:01 ZAFAR AHMED Septum if ON
  • 329522 07/03/10 01:33 N Toro ALWAYS coordinate the septum current with MCC! (**ALL SHIFT-TAKERS MUST READ**)
  • 329517 07/03/10 01:08 ZAFAR AHMED Cotrolled access and septum is off
  • 329516 07/03/10 00:39 Markowitz Swing shift summary
  • Friday

  • 329508 07/02/10 23:48 N Toro flashing light photo
  • 329497 07/02/10 22:45 adaq irregularly flashing light close to the target
  • 329495 07/02/10 22:15 Markowitz Beam back
  • 329494 07/02/10 22:00 Markowitz septum back on
  • 329493 07/02/10 21:09 Markowitz Hall survey in progress
  • 329492 07/02/10 20:42 N Toro septum FSD update
  • 329491 07/02/10 20:22 markowitz septum off
  • 329490 07/02/10 20:12 adaq magnet screens
  • 329489 07/02/10 20:00 J. Beacham Moved to empty target
  • 329488 07/02/10 19:55 adaq spot++ for BeO check
  • 329483 07/02/10 19:25 J. Beacham Move to BeO target
  • 329482 07/02/10 19:20 N Toro comments on corrector current
  • 329481 07/02/10 19:13 Markowitz More permanent magnet/ corrector coil commissioning
  • 329480 07/02/10 19:00 Markowitz More permanent magnet/ corrector coil commissioning
  • 329479 07/02/10 18:59 Markowitz More permanent magnet/ corrector coil commissioning
  • 329478 07/02/10 18:56 Markowitz First beam for permanent magnet/ corrector coil commissioning
  • 329477 07/02/10 18:20 Markowitz Right HRS Q2 and Q3 cycled
  • 329476 07/02/10 18:12 Ole Hansen Suggested VDC checkout
  • 329475 07/02/10 18:11 Jin Huang Optics Plots -> Replay Display
  • 329474 07/02/10 17:53 jpchen Target cooldown completed
  • 329473 07/02/10 17:49 J. Beacham Target cool-down status
  • 329472 07/02/10 17:34 Markowitz Cycling left HRS quads
  • 329471 07/02/10 17:04 riordan Run Plan for Tonight
  • 329468 07/02/10 16:57 jpchen switched from warm to cold retur
  • 329464 07/02/10 16:37 J. Beacham Target pressure and cool-down status
  • 329461 07/02/10 16:28 R. Michaels comments on coincidence measurements
  • 329459 07/02/10 16:17 Jin Huang Shift Summary : July 02 Day
  • 329458 07/02/10 16:07 K. Bartlett TA 67 HV line moved on right arm
  • 329456 07/02/10 15:32 R. Michaels restored cooling on ROC1 fastbus
  • 329453 07/02/10 15:05 jpchen Started to open J-T valve
  • 329447 07/02/10 14:10 Jin Huang Hall in Control Access
  • 329442 07/02/10 13:55 Jin Huang Target pressure history; Target cool down begin
  • 329441 07/02/10 13:52 jpchen Starting target (15K) cool-down
  • 329431 07/02/10 12:29 Jin Huang Done: all HRS magnets ->0A (except RHRS D1)
  • 329430 07/02/10 12:29 A. Camsonne Right arm gas cerenkov database updated
  • 329429 07/02/10 12:10 Folts checklist/tech
  • 329428 07/02/10 12:09 Jin Huang Start all HRS magnets ->0A (except RHRS D1)
  • 329427 07/02/10 11:54 Jin Huang Can not change RHRS D1 current
  • 329422 07/02/10 11:50 Fansler Target cooldown
  • 329417 07/02/10 11:02 fansler Quads
  • 329414 07/02/10 10:53 fansler right Q2
  • 329409 07/02/10 10:22 Bob / Sergei ADC datamap for R-HRS
  • 329406 07/02/10 09:49 R. Michaels ADCs all working now
  • 329390 07/02/10 07:47 Folts cycling magnets
  • 329371 07/02/10 05:49 A. Camsonne Front gate ADC still not working
  • 329370 07/02/10 05:48 A. Camsonne Front gate ADC still not working
  • 329358 07/02/10 00:29 D.Wang swapped out ROC1 slot 24 ADC board
  • Thursday

  • 329357 07/01/10 23:25 Bob / Bogdan R-HRS gate / timing adjustment
  • 329349 07/01/10 22:50 R. Michaels TDC channels for some R-HRS detectors
  • 329328 07/01/10 20:33 V. Sulkosky Right Q3 Should be Cycled to 1400 A
  • 329322 07/01/10 19:51 Folts chamber vacuum
  • 329320 07/01/10 19:34 LeRose Partial Shift Summary
  • 329318 07/01/10 19:14 K. Bartlett Installed 40 ns cable delays on trigger 1-5 on L-HRS
  • 329317 07/01/10 18:46 LeRose Q2/3 cycled
  • 329316 07/01/10 18:45 LeRose Right Dipole NMR lock
  • 329315 07/01/10 18:30 LeRose Right Arm Set
  • 329313 07/01/10 17:16 Neil G.W. Changes to RHS GC front end
  • 329312 07/01/10 17:05 Folts target chamber/tech
  • 329311 07/01/10 17:00 Bob / Kurtis new trigger T6, and timing adjustments
  • 329301 07/01/10 15:17 K. Bartlett/ Checked all NIM/ECL units
  • 329300 07/01/10 15:12 K. Bartlett/ Scaler Map Updated
  • 329291 07/01/10 14:13 R. Michaels TDC in roc1, slot 11 stopped again. Moved TDC
  • 329283 07/01/10 13:02 R. Michaels gate timing very precisely wrong
  • 329281 07/01/10 12:27 ejensen GC scaler rates
  • 329280 07/01/10 12:22 K. Bartlett R-HRS scalar map updated and scalar positions
  • 329276 07/01/10 12:05 R. Michaels timing adjustments / to-do
  • 329275 07/01/10 11:58 Bob / Neil TDC for trigger signals
  • 329269 07/01/10 11:44 folts target pix
  • 329264 07/01/10 10:46 Bob / Neil L-HRS gate timing check
  • 329250 07/01/10 08:15 kelleher techs: NMR not locking on Left Dipole
  • 329242 07/01/10 03:32 Jin Huang Optics Database Update
  • 329241 07/01/10 01:06 vmaxwell owl shift
  • 329240 07/01/10 00:56 vmaxwell owl shift summary